OpenStudy (anonymous):

I have to rewrite each function as a group of factors if possible. f(x)=x^2-16 g(x)=x^2+25 h(x)=(x+6)^2 j(x)=5x^2+10x-15

OpenStudy (anonymous):

\[ f(x)=x^2-16 \] \[g(x)=x^2+25\] \[ h(x)=(x+6)^2 \] \[j(x)=5x^2+10x-15 \]

OpenStudy (anonymous):

@Michele_Laino